OpenStudy (anonymous):

Find the values of x on the interval [0, π] where the tangent line to the graph of y = sin(x) cos(x) is horizontal. (Enter your answers as a comma-separated list.)

OpenStudy (anonymous):

y = sinxcosx dy/dx = cos^2x - sin^2x = cos2x For the tangent to be horizontal, dy/dx = 0 => cos2x = 0 2x = (2n+1)pi/2 x = (2n+1)pi/2 => x = pi/2 (only possible value since x lies in the interval [0,pi])